OpenStudy (anonymous):

Use logarithmic differentiation to find the derivative of the function. y=(sin(7x))^ln(x) y'=?

OpenStudy (anonymous):

ln(y)=In(x)[ln sin(7x)] (1/y)dy/dx= (1/x) ln sin(7x) +[7 cos (7x)/sin(7x)].lnx y'= y[ (1/x) ln sin (7x)+7 ln(x) .cot(7x)] y'=[sin(7x))^ln(x) ][ (1/x) ln sin (7x)+7 ln(x) .cot(7x)]